OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x^2 + 6x + 7 = 0

OpenStudy (anonymous):

x^2 + 6x + 9 = 9 - 7 ( x + 3)^2 = 2 x + 3 = sqrt(2)

OpenStudy (anonymous):

then

OpenStudy (anonymous):